OpenStudy (anonymous):

a^2 - 10ab +3b^2 complete factoring... is this prime ? because I cannot solve it

OpenStudy (anonymous):

\[a^2-10ab+3b^2=0\]

OpenStudy (anonymous):

quadratic formula is what you need

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yes, it is prime.

OpenStudy (anonymous):

\[-b \pm \sqrt{b^2-4ac}\over2a\]

OpenStudy (anonymous):

\[\pm\]

OpenStudy (anonymous):

that just means either or will yield the same answer.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

that will probably drive you insane though...be warned.

OpenStudy (anonymous):

Good luck with this one, I know it was fun for me :D

OpenStudy (anonymous):

You can factor it over the real numbers as \[\left(a-\sqrt{22} b-5 b\right) \left(a+\sqrt{22} b-5 b\right) \] I think there asking you to factor it over the integers. It is prime over the integers.

OpenStudy (anonymous):

Forget quadratic formula. That will probably be easier.

OpenStudy (anonymous):

YEA IT WAS MY LAST PROBLEM I PUT PRIME FOR SOME REASON I COUULDNT FIGURE IT OUT srry caps lok